OpenStudy (anonymous):

sine of x divided by one minus cosine of x + sine of x divided by one minus cosine of x = 2 csc x

OpenStudy (anonymous):

\[\frac{ \sin x }{ 1-\cos x }+\frac{ \sin x }{ 1-\cos x }\] check your statement.

OpenStudy (anonymous):

the second one, the denominator should be 1+cosx and that all equals 2csc x

OpenStudy (anonymous):

\[\frac{ \sin x }{ 1-\cos x }+\frac{ \sin x }{ 1+\cos x }=\sin x \left[ \frac{ 1+\cos x+1-\cos x }{ \left( 1-\cos x \right)\left( 1+\cos x \right) } \right]\] \[=\sin x \left[ \frac{ 2 }{ 1-\cos ^2x } \right]=\sin x * \frac{ 2 }{ \sin ^2x }=?\]